# SlimDiff: Training-Free, Activation-Guided Hands-free Slimming of Diffusion Models

### v1

@@ 1 Introduction
Diffusion models (DMs) (Rombach et al. (2022); Ramesh et al. (2022); Saharia et al. (2022)) have become the dominant paradigm in generative modeling, achieving remarkable performance. Their power, however, comes at a steep computational cost: every sample requires hundreds of denoising iterations, each iteration invoking a billion-parameter U-Net architecture (Dhariwal & Nichol (2021)). The sequential reliance on such high-dimensional operators leads to substantial latency, memory, and energy demands, making real-time or resource-constrained deployment prohibitive.
@@
Structural slimming offers a direct avenue for reducing both parameters and MACs (Shen et al. (2025)), yet applying it to DMs exposes fundamental challenges that prior work has largely overlooked. Attention (Vaswani et al. (2017)), a primary building block of the diffusion U-Net, illustrates this difficulty: pruning individual weights neglects the coupled nature of effective computations, which arise from products like query-key interactions ($\mathcal{QK}$), value-output couplings ($\mathcal{VO}$), and feedforward projections ($\mathcal{FFN}$). Since the rank of a product is bounded by the smallest rank among its factors (Kolter (2007)), maximal compression can be attained only when these products are treated as functional units (Lin et al. (2024)). As shown in App A.6, DM weights are largely high-rank with heavy-tailed spectra, and truncation errors accumulate through sequential denoising. Effective compression is achieved when the weight structure aligns with activation correlations, which define the active subspaces during denoising. Thus, optimal compression requires data awareness (Lin et al. (2024)). Ignoring these structural dependencies leads to subpar pruning decisions.
@@
The second issue is that compressibility in DMs is inherently timestep-dependent. Activation correlations reveal a far richer low-rank structure than weights alone, but the their covariance evolves dramatically over the denoising trajectory (Wang et al. (2024)). The activation distribution differs not only across timesteps but also across functional modules, each interacting with the weights in distinct ways. Approximating weights with a static, timestep-agnostic basis therefore collapses this evolving geometry and leads to poor preservation of fine details (Yao et al. (2024b)).
@@
A third issue is error propagation. In diffusion, pruning errors are not local; distortions introduced at one step are passed through every subsequent denoising update. Small deviations in early layers compound multiplicatively, creating irreversible degradation (Zeng et al. (2025)). Existing methods allocate sparsity myopically, without accounting for this sequential amplification, and therefore, rely on costly fine-tuning or retraining with large datasets to recover lost performance. Such dependence on retraining undermines the very motivation for slimming. (Zhang et al. (2024a))
@@
Finally, data-aware slimming requires collecting activations across timesteps and prompts (Lin et al. (2024)), and exhaustive sampling is computationally infeasible. Since naïve calibration over thousands of prompts is cumbersome, a principled strategy is needed to select a compact yet representative subset of prompts that spans the relevant activation subspace (Nguyen & He (2025)).
@@
Prior works ( Zhang et al. (2024b); Gao et al. (2024)) exemplify these limitations: although they reduce parameter counts, they remain functional module-agnostic and rely on finetuning or distillation on a large dataset to correct for timestep-dependent distortions and error propagation. Other approaches, such as  (Lu et al. (2022)) and ( Chen et al. (2025)), design compact models from scratch but at the cost of prohibitively expensive retraining. Alternatively, a parallel line of work orthogonally looks at inference-time accelerations such as (Wang et al. (2024); Bolya & Hoffman (2023); Fang et al. (2023)) that reduce computation by truncating denoising timesteps or merging tokens at run-time. These methods incur runtime overhead at every invocation and yield savings that fluctuate across runs - making both the effective cost and the achievable compression level unpredictable. Other methods explore quantization (Li et al. (2023); Zeng et al. (2025)) strategies to accelerate inference, which can be used in parallel with our structural slimming method.
@@
  • Comprehensive validation: We evaluate SlimDiff on MS-COCO (Lin et al. (2014)), LAION Aesthetics (Schuhmann et al. (2022)), ImageReward (Xu et al. (2023)), and PartiPrompts (Yu et al. (2023)) across DMs SDv1.5 and SDv1.4 (Rombach & Esser (2022b, a)). SlimDiff reduces $\sim 100$M parameters, reduces FLOPs by $22\%$, and speeds up inference by $35\%$ while preserving quality. We also confirm robustness via human preference scoring on HPS v2.1( Wu et al. (2023)), ImageReward, and Pic-a-Pic v1( Kirstain et al. (2023)).
@@ 2 Methodology
$\min_{\hat{\Theta}}\;\mathcal{L}_{\text{qual}}(\hat{\Theta})\quad\text{s.t.}\quad\mathrm{params}(\hat{\Theta})\leq B\vskip-7.0pt$ (1)
SlimDiff addresses this challenge through structural, activation-guided compression. Instead of pruning or factorizing individual matrices, it operates on functional weight groups ($\mathcal{QK}$, $\mathcal{VO}$, $\mathcal{FFN}$) and aligns compression with activation statistics across denoising steps. Figure 1 outlines four key components: (i) Spectral Influence Scoring: For each module, we quantify alignment with dominant activation directions to measure its relative importance (Sec. 2.1). (ii) Semantic Calibration Dataset (SlimSet): From the full prompt pool $\mathcal{P}$, we select a compact subset $\mathcal{S}$ that spans activations across timesteps, using geometric-median clustering with furthest-point sampling (FPS) in embedding space ( Nguyen & He (2025)) (Sec. 2.2). (iii) Timestep-aware Correlation Modeling: Using SlimSet activations, we compute per-timestep correlations for $\mathcal{QK}$, $\mathcal{VO}$, and $\mathcal{FFN}$ modules. These are aggregated into fidelity-weighted mixtures, assigning greater weight to timesteps most salient to output quality (Sec. 2.3). (iv) Module-Aligned Data-Aware Compression (MADAC) and Rank Allocation: Aggregated correlations drive a modular compression objective with tailored decompositions for $\mathcal{QK}$, $\mathcal{VO}$, and $\mathcal{FFN}$ blocks (Sec. 2.4). Influence scores then guide automatic per-layer rank selection under a user-specified parameter budget (Sec. 2.5). Together, these stages yield a closed-form, training-free pipeline that ‘slims’ diffusion models without performance degradation, delivering substantial reductions in parameters and FLOPs without retraining.
@@ 2.1 Anchoring Metric: Spectral Influence Score
At the core of SlimDiff is an anchoring metric we call the Spectral Influence Score, which quantifies how strongly each module aligns with the anisotropy of its input activations. This score serves as the foundation for both timestep-aware correlation accumulation and rank allocation, ensuring that compression decisions remain faithful to the evolving geometry of diffusion activations.
@@
$\mathcal{I}_{l,t}=\mathrm{TRQ}_{l,t}(\mathcal{W}_{l})=\frac{\mathcal{tr}(\mathcal{W}_{l}^{\top}\widetilde{\mathcal{C}}_{l,t}\mathcal{W}_{l})}{\|\mathcal{W}_{l}\|_{F}^{2}\,\mathcal{tr}(\widetilde{\mathcal{C}}_{l,t})}=\frac{\|\,\widetilde{\mathcal{C}_{l,t}}^{1/2}\mathcal{W}_{l}\,\|_{F}^{2}}{\|\mathcal{W}_{l}\|_{F}^{2}\,\mathrm{tr}(\widetilde{\mathcal{C}}_{l,t})}\vskip-4.0pt$ (2)
This formulation is variance-invariant: normalizing by $\mathrm{tr}(\widetilde{\mathcal{C}}_{l,t})$ cancels stepwise scaling, and scale-invariant, as normalizing by $|\mathcal{W}_{l}|_{F}^{2}$ removes dependence on parameter norms. Importantly, it isolates directional alignment: in the eigenbasis of $\widetilde{\mathcal{C}}_{l,t}$, the score evaluates how strongly $\mathcal{W}_{l}$ projects onto high-variance directions, focusing on anisotropy rather than raw energy. Aggregating TRQ across timesteps with a convex mixture (Sec. 2.3) yields a single spectral influence score per module $\mathcal{I}_{l,t}$, which anchors our pipeline by guiding both Data-Aware Compression and Global Rank Allocation.
@@ 2.2 SlimSet: Semantic-Aware Calibration Dataset Formation
Activation-guided compression requires estimating correlations $\Sigma_{l,t}=\mathbb{E}[X_{l,t}^{\top}X_{l,t}]$ across timesteps and modules. However, collecting these statistics over the full prompt corpus $\mathcal{P}$ is computationally expensive. We therefore introduce SlimSet, a semantic coreset $\mathcal{S}$ that preserves the statistical geometry of $\mathcal{P}$ while reducing calibration cost by more than $70\times$. Note that, SlimSet construction is lightweight, taking only a few minutes, and follows the semantic coreset selection strategy introduced in SCDP ( Nguyen & He (2025)).
Semantic embedding. Each prompt $p_{i}\in\mathcal{P}$ is embedded into $E_{i}\in\mathbb{R}^{d}$ using CLIP, providing a semantic space where distances reflect prompt similarity. We compute the geometric median $c$ of the embedding cloud, and assign each prompt a distinctiveness score $f_{i}=\|E_{i}-c\|_{2}$, so that prompts farther from the center capture more diverse semantics.
Bin allocation and sampling. To balance frequent and rare concepts, we stratify prompts into $B$ quantile bins of $\{f_{i}\}$ and assign each bin a quota $q_{b}$ proportional to corpus size. Within each bin, we apply farthest-point sampling (FPS): $\mathcal{S}_{b}=\arg\max_{|\mathcal{S}_{b}|=q_{b}}\;\min_{i\neq j\in\mathcal{S}_{b}}\big(1-\cos(E_{i},E_{j})\big)$, ensuring that selected prompts are well-spread and mutually diverse. We then perform cosine-based de-duplication to remove redundancy.
Resulting calibration set. The final SlimSet $\mathcal{S}$ retains only $J\ll|\mathcal{P}|$ prompts ($J=500$ vs. 3535k), yet yields activation covariances satisfying $\Sigma Cov_{l,t}^{\mathcal{S}}\approx\Sigma Cov_{l,t}^{\mathcal{P}}$ across layers $l$ and timesteps $t$. This compactness permits accurate estimation of module-level covariance structure at a fraction of the sampling cost, while preserving both semantic diversity and statistical fidelity. Fig. 3 and Sec. 4(i) detail the SlimSet size ablation and stability evaluation.
@@ 2.3 Timestep-Aware Correlation Modeling
DM activations evolve over the denoising trajectory, with correlation structure strongly tied to timestep. A single, timestep-agnostic covariance thus misrepresents the statistics: early steps are near-isotropic noise, while later ones show anisotropic, perceptually aligned variance (App. Sec A.5, Fig.  4). To capture this, we compute timestep-aware correlations per module and form a fidelity-weighted mixture emphasizing steps relevant to output quality. For activations $x_{l,t}\!\in\!\mathbb{R}^{d}$ collected at layer $l$ and timestep $t$ with $N_{t}$ samples, we form the second moment $\widehat{\mathcal{C}}_{l,t}=\tfrac{1}{N_{t}}\,\mathcal{X}_{l,t}^{\top}\mathcal{X}_{l,t}$. Next, to ensure numerical stability (especially with a few samples), we use a simple regularized estimate, denoted $\widetilde{\mathcal{C}}_{l,t}$, and then aggregate across timesteps using convex weights $w_{l,t}\!\geq\!0$, $\sum_{t}w_{l,t}=1$ chosen via the spectral influence score $\mathcal{I}_{l,t}$ (Sec. 2.1): $\bar{\mathcal{C}}_{l}=\sum_{t=1}^{T}w_{l,t}\,\widetilde{\mathcal{C}}_{l,t},\bar{\mathcal{R}}_{l}=\bar{\mathcal{C}}_{l}^{1/2}$. For cross-attention, the text features are time-invariant, so their statistics are computed once over SlimSet and reused.
Per-module variants. We apply spectral influence score $\mathcal{I}_{l,t}$ consistently across functional groups: (i) Self-attention (SA): $\mathcal{QK}$ logits use $\bar{\mathcal{R}}^{\text{sa}}$; $\mathcal{VO}$ maps use $\bar{\mathcal{R}}^{\text{sa}}$. (ii) Cross-attention (CA): Queries use step-mixtures $\bar{\mathcal{R}}^{q}$, while keys and values use cached text statistics $\mathcal{R}^{\text{text}}$. (iii) Feed-forward (FFN): Down-projection $\mathcal{W}_{d}$ is scored with FFN intermediate correlations $\bar{\mathcal{K}}^{\text{ffn}}$. We show the variation of $\mathcal{I}_{l,t}$ across layers, timesteps and functional modules in App. Sec  A.5, Fig  3.

### v2

@@ 1 Introduction
Diffusion models (DMs) (Rombach et al. (2022); Ramesh et al. (2022); Saharia et al. (2022)) have become the dominant paradigm in generative modeling, achieving remarkable performance. Their power, however, comes at a steep computational cost: every sample requires hundreds of denoising iterations, each iteration invoking a billion-parameter U-Net architecture (Dhariwal and Nichol (2021)). The sequential reliance on such high-dimensional operators leads to substantial latency, memory, and energy demands, making real-time or resource-constrained deployment prohibitive.
@@
Structural slimming offers a direct avenue for reducing both parameters and MACs (Shen et al. (2025)), yet applying it to DMs exposes fundamental challenges that prior work has largely overlooked. Attention (Vaswani et al. (2017)), a primary building block of the diffusion U-Net, illustrates this difficulty: compressing weight matrices in isolation neglects the coupled nature of effective computations, which arise from products like query-key interactions ($\mathcal{QK}$), value-output couplings ($\mathcal{VO}$), and feedforward projections ($\mathcal{FFN}$). Since the rank of a product is bounded by the smallest rank among its factors (Kolter (2007)), maximal compression can be attained only when these products are treated as functional units (Lin et al. (2024)). As shown in App A.6, DM weights are largely high-rank with heavy-tailed spectra, and truncation errors accumulate through sequential denoising. Effective compression is achieved when the weight structure aligns with activation correlations, which define the active subspaces during denoising. Thus, optimal compression requires data awareness (Lin et al. (2024)). Ignoring these structural dependencies leads to suboptimal compression decisions.
@@
The second issue is that compressibility in DMs is inherently timestep-dependent. Activation correlations reveal a far richer low-rank structure than weights alone, but the their covariance evolves dramatically over the denoising trajectory (Wang et al. (2024a)). The activation distribution differs not only across timesteps but also across functional modules, each interacting with the weights in distinct ways. Approximating weights with a static, timestep-agnostic basis therefore collapses this evolving geometry and leads to poor preservation of fine details (Yao and others (2024)).
@@
A third issue is error propagation. In diffusion, compression errors are not local; distortions introduced at one step are passed through every subsequent denoising update. Small deviations in early layers compound multiplicatively, creating irreversible degradation (Zeng et al. (2025)). Existing methods allocate sparsity myopically, without accounting for this sequential amplification, and therefore, rely on costly fine-tuning or retraining with large datasets to recover lost performance. Such dependence on retraining undermines the very motivation for slimming. (Zhang et al. (2024))
@@
Finally, data-aware slimming requires collecting activations across timesteps and prompts (Lin et al. (2024)), and exhaustive sampling is computationally infeasible. Since naïve calibration over thousands of prompts is cumbersome, a principled strategy is needed to select a compact yet representative subset of prompts that spans the relevant activation subspace (Nguyen and He (2025)).
@@
Prior works ( Zhang and others (2024); Kim et al. (2024)) exemplify these limitations: although they reduce parameter counts, they remain functional module-agnostic and rely on finetuning or distillation on a large dataset to correct for timestep-dependent distortions and error propagation. Other approaches, such as  (Lu et al. (2022)) and ( Chen et al. (2025)), design compact models from scratch but at the cost of prohibitively expensive retraining. Alternatively, a parallel line of work orthogonally looks at inference-time accelerations such as (Wang et al. (2024a); Bolya and Hoffman (2023); Fang et al. (2023)) that reduce computation by truncating denoising timesteps or merging tokens at run-time. These methods incur runtime overhead at every invocation and yield savings that fluctuate across runs - making both the effective cost and the achievable compression level unpredictable. Other methods explore quantization (Li et al. (2023); Zeng et al. (2025)) strategies to accelerate inference, which can be used in parallel with our structural slimming method.
@@
  • Comprehensive validation: We evaluate SlimDiff on MS-COCO (Lin et al. (2014)), LAION Aesthetics (Schuhmann et al. (2022)), ImageReward (Xu et al. (2023)), and PartiPrompts (Yu et al. (2023)) across DMs SDv1.5 and SDv1.4 (Rombach and Esser (2022b, a)). SlimDiff reduces $\sim 100$M parameters, reduces FLOPs by $22\%$, and speeds up inference by $35\%$ while preserving quality. We also confirm robustness via human preference scoring on HPS v2.1( Wu et al. (2023)), ImageReward, and Pic-a-Pic v1( Kirstain et al. (2023)).
@@ 2 Methodology
$\min_{\hat{\Theta}}\;\mathcal{L}_{\text{qual}}(\hat{\Theta})\quad\text{s.t.}\quad\mathrm{params}(\hat{\Theta})\leq B$ (1)
where $\mathcal{L}_{\text{qual}}(\hat{\Theta})=\mathbb{E}_{p,t}\left[\|f_{\Theta}(x_{t},p)-f_{\hat{\Theta}}(x_{t},p)\|^{2}\right]$ measures the expected reconstruction error between the original and compressed models across prompts $p$ and timesteps $t$. Rather than optimizing this intractable objective directly, SlimDiff decomposes it into per-module surrogate objectives (Eq. 5) that admit closed-form solutions. These closed-form surrogates operate at the level of functional weight groups: self-attention query–key pairs ($\mathcal{QK}$), value–output pairs ($\mathcal{VO}$), and feed-forward blocks ($\mathcal{FFN}$), rather than individual matrices. For each group, SlimDiff uses activation correlations to construct a activation-guided low-rank approximation that minimizes $\mathcal{L}_{\text{qual}}$ locally under a rank constraint. Our pipeline in figure 1 outlines four key components: (i) Spectral Influence Scoring: For each module, we quantify alignment with dominant activation directions to measure its relative importance (Sec. 2.1). (ii) Semantic Calibration Dataset (SlimSet): From the full prompt pool $\mathcal{P}$, we select a compact subset $\mathcal{S}$ that spans activations across timesteps, using geometric-median clustering with furthest-point sampling (FPS) in embedding space ( Nguyen and He (2025)) (Sec. 2.2). (iii) Timestep-aware Correlation Modeling: Using SlimSet activations, we compute per-timestep correlations for $\mathcal{QK}$, $\mathcal{VO}$, and $\mathcal{FFN}$ modules. These are aggregated into fidelity-weighted mixtures, assigning greater weight to timesteps most salient to output quality (Sec. 2.3). (iv) Module-Aligned Data-Aware Compression (MADAC) and Rank Allocation: Aggregated correlations drive a modular compression objective with tailored decompositions for $\mathcal{QK}$, $\mathcal{VO}$, and $\mathcal{FFN}$ blocks (Sec. 2.4). Influence scores then guide a propagation-aware per-layer rank selection under a user-specified parameter budget (Sec. 2.5), assigning higher ranks to modules whose errors accumulate most strongly over the denoising trajectory. Together, these stages yield a closed-form, training-free pipeline that ‘slims’ diffusion models without performance degradation, delivering substantial reductions in parameters and FLOPs without retraining.
@@ 2.1 Anchoring Metric: Spectral Influence Score
At the core of SlimDiff is an anchoring metric, we call the Spectral Influence Score, which quantifies how strongly each module aligns with the anisotropy of its input activations. This score serves as the foundation for both timestep-aware correlation accumulation and rank allocation, ensuring that compression decisions remain faithful to the evolving geometry of diffusion activations.
@@
$\mathcal{I}_{l,t}=\mathrm{TRQ}_{l,t}(\mathcal{W}_{l})=\frac{\text{tr}(\mathcal{W}_{l}^{\top}\widetilde{\mathcal{C}}_{l,t}\mathcal{W}_{l})}{\|\mathcal{W}_{l}\|_{F}^{2}\,\text{tr}(\widetilde{\mathcal{C}}_{l,t})}=\frac{\|\,\widetilde{\mathcal{C}_{l,t}}^{1/2}\mathcal{W}_{l}\,\|_{F}^{2}}{\|\mathcal{W}_{l}\|_{F}^{2}\,\text{tr}(\widetilde{\mathcal{C}}_{l,t})}\vskip-4.0pt$ (2)
This formulation is variance-invariant: normalizing by $\mathrm{tr}(\widetilde{\mathcal{C}}_{l,t})$ cancels stepwise scaling, and scale-invariant, as normalizing by $|\mathcal{W}_{l}|_{F}^{2}$ removes dependence on parameter norms. Importantly, it isolates directional alignment: in the eigenbasis of $\widetilde{\mathcal{C}}_{l,t}$, the score evaluates how strongly $\mathcal{W}_{l}$ projects onto high-variance directions, focusing on anisotropy rather than raw energy. Aggregating TRQ across timesteps with a convex mixture (Sec. 2.3) yields a single spectral influence score per module $\mathcal{I}_{l}$, which we use to construct a single data-aware correlation per module and to anchor the overall compression objective. These scores drive a propagation-aware rank allocation, assigning higher ranks to modules whose anisotropic directions contribute most to error accumulation along the denoising trajectory.
@@ 2.2 SlimSet: Semantic-Aware Calibration Dataset Formation
Activation-guided compression requires estimating correlations $\Sigma_{l,t}=\mathbb{E}[X_{l,t}^{\top}X_{l,t}]$ across timesteps and modules. However, collecting these statistics over the full prompt corpus $\mathcal{P}$ is computationally expensive. We therefore introduce SlimSet, a semantic coreset $\mathcal{S}$ that preserves the statistical geometry of $\mathcal{P}$ while reducing calibration cost by more than $70\times$. Note that, SlimSet construction is lightweight, taking only a few minutes, and follows the semantic coreset selection strategy introduced in SCDP ( Nguyen and He (2025)).
Semantic embedding. Each prompt $p_{i}\in\mathcal{P}$ is embedded into $E_{i}\in\mathbb{R}^{d}$ using CLIP, providing a semantic space where distances reflect prompt similarity. We compute the geometric median $c$ of the embedding cloud, and assign each prompt a distinctiveness score $f_{i}=\|E_{i}-c\|_{2}$ (also denoted as FD). Large $f_{i}$ (far from median $c$) captures rare semantics, small $f_{i}$ (near $c$) captures common semantics.
Figure 2: SlimSet coverage. Distribution of distinctiveness scores for LAION-212K (grey) and SlimSet with J=500J{=}500 (blue). Quantile-based binning with proportional allocation ensures SlimSet spans the entire corpus range.
Bin allocation and sampling. To balance frequent and rare concepts, we stratify prompts into $B$ quantile bins of $\{f_{i}\}$ and assign each bin a quota $q_{b}$ proportional to corpus size. Within each bin, we apply farthest-point sampling (FPS):
$\mathcal{S}_{b}=\arg\max_{|\mathcal{S}_{b}|=q_{b}}\;\min_{i\neq j\in\mathcal{S}_{b}}\bigl(1-\cos(E_{i},E_{j})\bigr),$
, ensuring selected prompts are well-spread and mutually diverse. We then perform cosine-based de-duplication to remove redundancy. This two-stage strategy: quantile stratification followed by within-bin FPS, guarantees coverage of both central (low-$f$) and tail (high-$f$) semantics.
Resulting calibration set. The final SlimSet $\mathcal{S}$ retains only $J\ll|\mathcal{P}|$ prompts ($J=500$ vs. 3535k), yet yields activation covariances satisfying $\Sigma Cov_{l,t}^{\mathcal{S}}\approx\Sigma Cov_{l,t}^{\mathcal{P}}$ across layers $l$ and timesteps $t$. This compactness permits accurate estimation of module-level covariance structure at a fraction of the sampling cost, while preserving both semantic diversity and statistical fidelity. Fig. 3 and Sec. 4(i) detail the SlimSet size ablation and stability evaluation. Appx.7 provides per-dataset coverage analysis and qualitative examples from all quantile bins.
@@ 2.3 Timestep-Aware Correlation Modeling
DM activations evolve over the denoising trajectory, with correlation structure strongly tied to timestep. A single, timestep-agnostic covariance thus misrepresents the statistics: early steps are near-isotropic noise, while later ones show anisotropic, perceptually aligned variance (App. Sec A.5, Fig.  5). To capture this, we compute timestep-aware correlations per module and form a fidelity-weighted mixture emphasizing steps relevant to output quality.
Let $x_{l,t}^{(i)}\in\mathbb{R}^{d}$ denote the $i$-th activation sample (corresponding to the $i$-th prompt) at layer $l$ and timestep $t$. Let $\mathcal{X}_{l,t}\in\mathbb{R}^{N_{t}\times d}$ denote the data matrix stacking all $N_{t}$ spatial samples across prompts. We compute the empirical second moment as:
$\widehat{\mathcal{C}}_{l,t}=\tfrac{1}{N_{t}}\,\mathcal{X}_{l,t}^{\top}\mathcal{X}_{l,t}=\tfrac{1}{N_{t}}\sum_{i=1}^{N_{t}}x_{l,t}^{(i)}(x_{l,t}^{(i)})^{\top}\in\mathbb{R}^{d\times d}$ (3)
To ensure numerical stability and invertibility (especially with limited prompts), we use a regularized estimate: $\widetilde{\mathcal{C}}_{l,t}=\widehat{\mathcal{C}}_{l,t}+\epsilon\mathbf{I}_{d}$, $\epsilon=10^{-6}$. This regularized correlation $\widetilde{\mathcal{C}}_{l,t}$ is used in the spectral influence score (Eq. 2). Finally, we aggregate across timesteps using convex weights $w_{l,t}\geq 0$, $\sum_{t}w_{l,t}=1$, derived from the spectral influence scores :
$\bar{\mathcal{C}}_{l}=\sum_{t=1}^{T}w_{l,t}\,\widetilde{\mathcal{C}}_{l,t},\quad\bar{\mathcal{R}}_{l}=\bar{\mathcal{C}}_{l}^{1/2}$ (4)
For cross-attention, the text features are time-invariant, so their correlation statistics are computed once over SlimSet and reused.
Per-module variants. We apply spectral influence score $\mathcal{I}_{l,t}$ consistently across functional groups: (i) Self-attention (SA): $\mathcal{QK}$ logits use $\bar{\mathcal{R}}^{\text{sa}}$; $\mathcal{VO}$ maps use $\bar{\mathcal{R}}^{\text{sa}}$. (ii) Cross-attention (CA): Queries use step-mixtures $\bar{\mathcal{R}}^{q}$, while keys and values use cached text statistics $\mathcal{R}^{\text{text}}$. (iii) Feed-forward (FFN): Down-projection $\mathcal{W}_{d}$ is scored with FFN intermediate correlations $\bar{\mathcal{K}}^{\text{ffn}}$. We show the variation of $\mathcal{I}_{l,t}$ across layers, timesteps and functional modules in App. Sec  A.5, Fig  4.